OpenStudy (anonymous):

Medalsss Solve for x: 8x + 17 = 5(x + 1) x = -8 x = 8 x = -4 x = 4

OpenStudy (cj49):

eg if 4x+8=2(x+6) 4x+8=2x+12 4x-2x=12-8 2x=4 x=2

OpenStudy (anonymous):

Is it -4

OpenStudy (cj49):

3x=-12 x=-4

OpenStudy (anonymous):

Can you help me with another one.